OpenStudy (anonymous):

Sarah investment in stock grew 22% to $366. how much did she invest?

OpenStudy (lgbasallote):

x + 0.22x = 366 1.22x = 366 x = $300

OpenStudy (anonymous):

A 108-inch broad is cut into two pieces. one piece is two times the length of the other. Find the length of the two pieces. The short piece_____ The long piece_____

OpenStudy (lgbasallote):

let x = one piece 2x = other piece x + 2x = 108 3x = 108 x = 36 2x = 64 so the shorter piece is 36...longer piece is 64

OpenStudy (anonymous):

A=yv+ty for v

OpenStudy (lgbasallote):

A = y(v+t) A/y = v + t v = (A/y) - t

OpenStudy (anonymous):

The perimeter of a basketball court is 90 meters and the length is 6 meters longer than twice the width . What are the length and width

OpenStudy (lgbasallote):

perimeter = l + l + w + w l = 6 + 2w 90 = (6+2w) + (6+ 2w) + w + w 90 = 12 + 4w + w +w 90 = 12 + 6w 78 = 6w w = 13 l = 32